A car undergoes a constant acceleration of 6.0 m/s2

starting from rest. In the first second, it travels ____?
1
Expert's answer
2020-12-17T07:27:39-0500

Answer

Using second equation of motion

"s=ut+1\/2at^2"

u=0 and the=1sec

 In the first second, it travels

As

"s=0+0.5\times6\times1=3m"